OpenStudy (anonymous):

if cosx=1/3 and 0<=beta<=90 what is the value of sin?

OpenStudy (anonymous):

Here are a couple hints: 1. If \(\cos\beta = 1/3\) and \(0\leq \beta\leq 90\), then \(\sin\beta\geq 0\). 2. Recall that \(\sin^2\beta+\cos^2\beta = 1\). Can you take things from here?